# Phase transitions from linear to nonlinear information processing in neural networks

### v1

@@ I Introduction
Reservoir computing is a computational framework that harnesses the information processing capacity of high-dimensional nonlinear dynamical systems, known as reservoirs, to efficiently process time series data Jaeger-09 ; Tanaka-19 ; Nakajima-21 ; Cucchi-22 . Unlike traditional neural networks where all layers are trained, reservoir computing keeps the reservoir structure fixed, training only a simple readout layer. The reservoir acts as a nonlinear mapping that projects time series data into a high-dimensional state space, where complex patterns and temporal dependencies are more easily separable. This approach simplifies the training process and has been effectively applied to tasks such as time series prediction Pathak-17 ; Lu-17 ; Pathak-18 ; Vlachas-20 ; Gauthier-21 ; Zhang-21 ; Kobayashi-21 ; Rohm-21 ; Srinivasan-22 ; Storm-22 ; Tanaka-22 ; Zhang-23 and speech recognition Verstraeten-05 ; Skowronski-07 ; Rodan-11 , using both virtual reservoirs like recurrent neural networks and physical reservoirs like optical and mechanical systems Appeltant-11 ; Nakajima-15 ; Larger-17 ; Du-17 ; Dion-18 ; Moon-19 ; Rafayelyan-20 ; Yada-21 ; Maksymov-22 ; Toprasertpong-22 .
@@
In the following, we demonstrate a phase transition from linear to nonlinear information processing in echo state networks, which are a standard paradigm in reservoir computing using recurrent neural networks. The reservoir consists of randomly connected nodes perturbed by a small amount of noise. The output is computed as a linear combination of the reservoir states, where weights are trained via standard linear regression to minimize the distance between the output and the target data. The information processing capacity is measured by the correlation coefficient between the output and the target. The nonlinearity of the system is controlled by the standard deviation of the input weights. When this value is small, the reservoir states remain small and are minimally influenced by the nonlinearity of the activation function. By examining the information processing capacity as a function of the input weight standard deviation, we observe a rapid increase in the capacity from the linear to the nonlinear regime at a critical value, with the transition becoming sharper as the number of nodes increases [see Fig. 2(a)-(d)]. This suggests the possibility of singular behavior in the information processing capacity in the limit of an infinite number of nodes. Additionally, we find that as the noise intensity decreases, the standard deviation of the input weights at which the transition occurs approaches zero, implying that even an infinitesimally weak nonlinearity can drive the system into the nonlinear information processing regime.
@@
This paper is organized as follows. In Sec. II, we introduce a general framework for defining the information processing capacity of dynamical systems. By adopting the reservoir computing perspective, the capacity is formulated as a measure of the system’s ability to reproduce a given input-output relationship. In Sec. III, we describe the echo state network model perturbed by a small amount of noise. We detail the initialization procedure for the recurrent and input weight matrices, and explain the method used to calculate the Lyapunov exponent of the network. In Sec. IV, we present our main results on the information processing capacity. The phase diagram shown in Fig. 2 highlights the existence of distinct linear and nonlinear information processing regimes. We further analyze how the capacity depends on the noise intensity and network size. Sec. V concludes the paper with a summary and outlook for future research. Appendix A provides the detailed procedure for training the echo state network. In Appendix B, we examine how the main results are modified when ridge regression is used instead of standard linear regression. Appendix C discusses the behavior of the information processing capacity for the NARMA and delay tasks. Finally, in Appendix D, we address the impact of round-off errors in numerical calculations, which become significant when both the noise intensity and input weight magnitude are extremely small.
@@ III Echo state network
The input weights $\mathbf{v}=(v_{1},...,v_{N})$ are drawn from a uniform distribution on $[-\sqrt{3}\sigma_{\text{in}},\sqrt{3}\sigma_{\text{in}}]$, where $\sigma_{\text{in}}$ is the standard deviation that controls the strength of the input. In a nonlinear network with the activation function $h(\mathbf{x})=\tanh(\mathbf{x})$, the input standard deviation $\sigma_{\text{in}}$ governs the degree of the nonlinearity: if $\sigma_{\text{in}}\ll 1$, then $|x_{i}(t)|\ll 1$, making $\tanh(\mathbf{x})\simeq\mathbf{x}$ effectively linear. The elements of the recurrent weight matrix $W_{ij}$ are first drawn from a uniform distribution on $[-1,1]$ with probability $p$, and set to zero with probability $1-p$. The matrix $W$ is then rescaled so that its spectral radius equals a prescribed value $\rho$. Throughout our calculations, we fix the connectivity factor $p=0.1$.
@@ IV.1 Phase diagram
We identify four phases (1)-(4), as highlighted in Fig. 2(b). In phase (1), characterized by small $\sigma_{\text{in}}$ and $\rho$, the system effectively operates as a linear network, so its capacity is zero for the XOR task. In phase (2), nonlinearity becomes significant, and the network achieves a high capacity. In phase (3), where $\rho$ is large, the network enters a chaotic regime with a positive Lyapunov exponent, and the capacity drops to zero. In phase (4), large $\sigma_{\text{in}}$ drives the state $x_{i}(t)$ to saturate at ±$\pm 1$ under the $\tanh$ activation, preventing effective nonlinear transformations. Our primary focus is the boundary between phases (1) and (2). Remarkably, as the number of nodes $N$ increases, this phase boundary becomes sharper, as also seen in Fig. 2(d). This observation suggests that the capacity $r^{2}$ exhibits a discontinuous jump in the limit $N\to\infty$, as illustrated in Fig. 1.
Figures 2(e)-(h) show the fluctuation $\sigma(r^{2})$ of the capacity across different network realizations. A pronounced peak appears at the boundary between phases (1) and (2). This is reminiscent of the divergence of order-parameter fluctuations observed at second-order phase transitions if one considers the capacity $r^{2}$ as the order parameter. However, there are notable differences from the conventional phase transition picture. First, our transition appears to be first-order (discontinuous), rather than second-order (continuous). Second, the height of the peak of $\sigma(r^{2})$ does not diverge in the limit $N\to\infty$. Finally, the critical input standard deviation $\sigma_{\text{in},0}$ at which the transition occurs gradually decreases as $N$ increases.
@@ IV.2 Scaling for the noise intensity
Let us consider how the noise $\bm{\xi}$ affects the capacity $r^{2}$. Figure 3 shows $r^{2}$ as a function of $\sigma_{\text{in}}$ for various noise intensities $\sigma_{\xi}$. As expected, large noise diminishes the capacity. Notably, in Fig. 3(b), when $r^{2}$ is plotted against the rescaled input standard derivation $\sigma_{\text{in}}/\sigma_{\xi}^{1/3}$, the curves corresponding to various $\sigma_{\xi}$ collapse onto a single master curve. We thus obtain a scaling law for the critical input standard deviation,
@@
Figure 3: Information processing capacity $r^{2}$ as a function of the input standard deviation $\sigma_{\text{in}}$ for various noise intensities $\sigma_{\xi}=10^{-10},10^{-9},10^{-8},10^{-7},10^{-6}$. The network has $N=1000$ nodes, and the spectral radius of $W$ is $\rho=0.8$. The XOR task uses a delay of $k=10$. In panel (b), $r^{2}$ is plotted against the rescaled input standard deviation $\sigma_{\text{in}}/\sigma_{\xi}^{1/3}$. The collapse of the curves for different $\sigma_{\xi}$ onto a single master curve is clearly observed.
@@
where $\beta$ is the ridge parameter. As discussed in Appendix B, $\beta$ corresponds to an effective noise standard deviation $\sigma_{\xi}\sim\beta^{1/2}$. Consequently, in the ridge regression case, the appropriate rescaling of the input standard deviation is given by $\sigma_{\text{in}}/\beta^{1/6}$, as demonstrated in Fig. 6.
@@ IV.3 Scaling for the network size
We define the rescaled critical point $\tilde{\sigma}_{\text{in},0}$ as the value of $\tilde{\sigma}_{\text{in}}$ at which the derivative $dr^{2}/d\tilde{\sigma}_{\text{in}}$ is maximized, denoting its maximum by $\Delta r^{2}_{\text{max}}$:
$\tilde{\sigma}_{\text{in},0}=\mathop{\rm argmax}\limits_{\tilde{\sigma}_{\text{in}}}\frac{dr^{2}(\tilde{\sigma}_{\text{in}})}{d\tilde{\sigma}_{\text{in}}},\quad\Delta r^{2}_{\text{max}}=\frac{dr^{2}(\tilde{\sigma}_{\text{in},0})}{d\tilde{\sigma}_{\text{in}}}.$ (16)
@@
To reduce the fluctuations in $\tilde{\sigma}_{\text{in},0}$ and $\Delta r^{2}_{\text{max}}$, we fit $r^{2}$ as a function of $\tilde{\sigma}_{\text{in}}$ using a logarithmic logistic function of the form,
$f(x)=\frac{a}{1+e^{-b(\ln x-c)}},$ (17)
with fitting parameters $a$, $b$, and $c$ [see Fig. 4(a)]. We then obtain $\tilde{\sigma}_{\text{in},0}$ and $\Delta r^{2}_{\text{max}}$ directly from these fitted parameters:
$\tilde{\sigma}_{\text{in},0}$ $=e^{c}\left(\frac{b-1}{b+1}\right)^{1/b},$ (18)
@@
Figure 4: Network size scaling of the transition point. (a) Fitting $r^{2}$ versus $\tilde{\sigma}_{\text{in}}$ using a logarithmic logistic function. The dashed line indicates $\tilde{\sigma}_{\text{in},0}$, at which $dr^{2}/d\tilde{\sigma}_{\text{in}}$ attains its maximum $\Delta r^{2}_{\text{max}}$. (b) Network size dependence of the maximum derivative $\Delta r^{2}_{\text{max}}$ at the transition point in a double-log plot. The dashed lines represent $\Delta r^{2}_{\text{max}}\propto N^{\alpha}$. (c) Network size dependence of the transition point $\tilde{\sigma}_{\text{in},0}$ in a double-log plot. The dashed lines represent $\tilde{\sigma}_{\text{in},0}\propto N^{-\beta}$. (d) Exponents $\alpha$ and $\beta$ for different delay values $k$. The power-law fits for $\tilde{\sigma}_{\text{in},0}$ and $\Delta r^{2}_{\text{max}}$ are performed for $N\geq 2200$. The error bars indicate the standard errors from the fitting. The fact that $\alpha>\beta$ implies that the transition sharpens faster than $\tilde{\sigma}_{\text{in},0}$ vanishes.
Figures 4(b) and (c) show the size dependence of $\Delta r^{2}_{\text{max}}$ and $\tilde{\sigma}_{\text{in},0}$ in a double-log plot for different delay values $k$ in the XOR task. For sufficiently large $N$, we find that the sharpness of the transition scales as $\Delta r^{2}_{\text{max}}\propto N^{\alpha}$, while its location scales as $\tilde{\sigma}_{\text{in},0}\propto N^{-\beta}$, with some exponents $\alpha$ and $\beta$. Figure 4(d) shows the exponents $\alpha$ and $\beta$ as functions of $k$. The power-law fits for $\tilde{\sigma}_{\text{in},0}$ and $\Delta r^{2}_{\text{max}}$ are performed for $N\geq 2200$. As $k$ increases, the XOR task demands longer memory, making it more difficult. We can observe that both $\alpha$ and $\beta$ increase with $k$, while it is unclear to what values they converge as $k\to\infty$. Crucially, $\alpha>\beta$, which implies that the transition sharpens more quickly than $\tilde{\sigma}_{\text{in},0}$ goes to zero. This implies that, if one plots $r^{2}$ against $\tilde{\sigma}_{\text{in}}N^{\beta}$, a discontinuous transition may emerge in the limit $N\to\infty$, as illustrated in Fig. 1.

### v2

@@ I Introduction
Reservoir computing is a computational framework that harnesses the information processing capacity of high-dimensional nonlinear dynamical systems, known as reservoirs, to efficiently process time series data Jaeger-09 ; Tanaka-19 ; Nakajima-21 ; Cucchi-22 . Unlike traditional neural networks where all layers are trained, reservoir computing keeps the reservoir structure fixed, training only a simple readout layer. The reservoir acts as a nonlinear mapping that projects time series data into a high-dimensional state space, where complex patterns and temporal dependencies are more easily separable. This approach simplifies the training process and has been effectively applied to tasks such as time series prediction Pathak-17 ; Lu-17 ; Pathak-18 ; Vlachas-20 ; Gauthier-21 ; Zhang-21 ; Kobayashi-21 ; Rohm-21 ; Srinivasan-22 ; Storm-22 ; Tanaka-22 ; Zhang-23 and speech recognition Verstraeten-05 ; Skowronski-07 ; Rodan-11 , using both virtual reservoirs like recurrent neural networks and physical reservoirs like optical and mechanical systems Appeltant-11 ; Nakajima-15 ; Dion-18 ; Du-17 ; Moon-19 ; Toprasertpong-22 ; Yada-21 ; Maksymov-22 ; Larger-17 ; Rafayelyan-20 ; Wang-24 .
@@
In the following, we demonstrate a phase transition from linear to nonlinear information processing in echo state networks, which are a standard paradigm in reservoir computing using recurrent neural networks. The reservoir consists of randomly connected nodes perturbed by a small amount of noise. The output is computed as a linear combination of the reservoir states, where weights are trained via standard linear regression to minimize the distance between the output and the target data. The information processing capacity is measured by the correlation coefficient between the output and the target. The nonlinearity of the system is controlled by the standard deviation of the input weights. When this value is small, the reservoir states remain small and are minimally influenced by the nonlinearity of the activation function. By examining the information processing capacity as a function of the input weight standard deviation, we observe a rapid increase in the capacity from the linear to the nonlinear regime at a critical value, with the transition becoming sharper as the number of nodes increases [see Fig. 2(a)-(d)]. This suggests the possibility of singular behavior in the information processing capacity in the limit of an infinite number of nodes. Additionally, we find that as the noise intensity decreases, the standard deviation of the input weights at which the phase transition occurs approaches zero, implying that even an infinitesimally weak nonlinearity can drive the system into the nonlinear information processing regime.
Here, we clarify our use of the term “phase transition”. In this study, we use it in a broad sense to describe a collective phenomenon in which an observable exhibits non-analytic behavior as a function of a control parameter in the limit of infinite system size. This definition highlights the distinction between our transition and bifurcation phenomena such as order-to-chaos transitions, which can occur even in finite systems. In particular, our transition does not necessarily display features commonly associated with conventional thermodynamic phase transitions, such as the divergence of correlation length or correlation time.
@@
This paper is organized as follows. In Sec. II, we introduce a general framework for defining the information processing capacity of dynamical systems. By adopting the reservoir computing perspective, the capacity is formulated as a measure of the system’s ability to reproduce a given input-output relationship. In Sec. III, we describe the echo state network model perturbed by a small amount of noise. We detail the initialization procedure for the recurrent and input weight matrices, and explain the method used to calculate the Lyapunov exponent of the network. In Sec. IV, we present our main results on the information processing capacity. The phase diagram shown in Fig. 2 highlights the existence of distinct linear and nonlinear information processing regimes. We further analyze how the capacity depends on the noise intensity and network size. Sec. V concludes the paper with a summary and outlook for future research. Appendix A provides the detailed procedure for training the echo state network. Appendix B details a finite-size scaling analysis of the transition. In Appendix C, we examine how the main results are modified when ridge regression is used instead of standard linear regression. Appendix D discusses the noise dependence of the information processing capacity for different tasks. In Appendix E, we examine how the choice of activation function influences the noise dependence of the transition point. Finally, in Appendix F, we address the impact of round-off errors in numerical calculations, which become significant when both the noise intensity and input weight magnitude are extremely small.
@@ III Echo state network
The input weights $\mathbf{v}=(v_{1},...,v_{N})$ are drawn from a uniform distribution on $\sqrt{3}\sigma_{\text{in}},\sqrt{3}\sigma_{\text{in}}]$, where $\sigma_{\text{in}}$ is the standard deviation that controls the strength of the input. In a nonlinear network with the activation function $h(\mathbf{x})=\tanh(\mathbf{x})$, the input standard deviation $\sigma_{\text{in}}$ governs the degree of the nonlinearity: if $\sigma_{\text{in}}\ll 1$, then $|x_{i}(t)|\ll 1$, making $\tanh(\mathbf{x})\simeq\mathbf{x}$ effectively linear. The elements of the recurrent weight matrix $W_{ij}$ are first drawn from a uniform distribution on [1,1][-1,1] with probability $p$, and set to zero with probability 1p1-p. The matrix $W$ is then rescaled so that its spectral radius equals a prescribed value $\rho$. Throughout our calculations, we fix the connectivity factor $p=0.1$.
@@ IV.1 Phase diagram
We identify four phases (1)-(4), as highlighted in Fig. 2(b). In phase (1), characterized by small $\sigma_{\text{in}}$ and $\rho$, the system effectively operates as a linear network, so its capacity is zero for the XOR task. In phase (2), nonlinearity becomes significant, and the network achieves a high capacity. In phase (3), where $\rho$ is large, the network enters a chaotic regime with a positive Lyapunov exponent, and the capacity drops to zero. In phase (4), large $\sigma_{\text{in}}$ drives the state $x_{i}(t)$ to saturate at $\pm 1$ under the $\tanh$ activation, preventing effective nonlinear transformations. Our primary focus is the boundary between phases (1) and (2). Remarkably, as the number of nodes $N$ increases, this phase boundary becomes sharper, as also seen in Fig. 2(d). This observation suggests that the capacity $r^{2}$ exhibits a discontinuous jump in the limit $N\to\infty$, as illustrated in Fig. 1.
Figures 2(e)-(h) show the fluctuation $\sigma(r^{2})$ of the capacity across different network realizations. A pronounced peak appears at the boundary between phases (1) and (2). This is reminiscent of the divergence of order-parameter fluctuations observed at second-order phase transitions if one considers the capacity $r^{2}$ as the order parameter. However, there are notable differences from the conventional phase transition picture. First, our transition appears to be first-order (discontinuous), rather than second-order (continuous). Second, the height of the peak of $\sigma(r^{2})$ does not diverge in the limit $N\to\infty$. This bounded behavior arises from the fact that $r^{2}$ itself is confined to the interval $0\leq r^{2}\leq 1$, in contrast to extensive order parameters in traditional systems, such as magnetization in the Ising model, which scale with system size. Finally, the critical input standard deviation $\sigma_{\text{in},0}$ at which the transition occurs gradually decreases as $N$ increases.
@@ IV.2 Scaling for the noise intensity
Figure 3: Scaling of the transition point with respect to the noise intensity $\sigma_{\xi}$. (a) Information processing capacity $r^{2}$ as a function of the input standard deviation $\sigma_{\text{in}}$ for various noise intensities $\sigma_{\xi}=10^{-10}$, $10^{-9}$, $10^{-8}$, $10^{-7}$, $10^{-6}$. The network has $N=1000$ nodes, and the spectral radius of $W$ is $\rho=0.8$. The XOR task uses a delay of $k=10$. (b) $r^{2}$ plotted against the rescaled input standard deviation $\sigma_{\text{in}}/\sigma_{\xi}^{1/3}$. The collapse of the curves for different $\sigma_{\xi}$ onto a single master curve is clearly observed. (c) Example of fitting $r^{2}$ versus $\sigma_{\text{in}}$ for $\sigma_{\xi}=10^{-8}$ using a logarithmic logistic function. The dashed line indicates the transition point $\sigma_{\text{in},0}$, defined as the point where the derivative $dr^{2}/d\sigma_{\text{in}}$ attains its maximum. (d) $\sigma_{\text{in},0}$ as a function of the noise intensity $\sigma_{\xi}$ for different spectral radii $\rho=0.4$, $0.6$, $0.8$. The dashed lines show power-law fits of the form $\sigma_{\text{in},0}\propto\sigma_{\xi}^{\eta}$, with $\eta=0.340$ for $\rho=0.4$, $\eta=0.334$ for $\rho=0.6$, and $\eta=0.333$ for $\rho=0.8$.
Let us consider how the noise affects the capacity $r^{2}$. Figure 3(a) shows $r^{2}$ as a function of $\sigma_{\text{in}}$ for various noise intensities $\sigma_{\xi}$. As expected, large noise diminishes the capacity. Notably, in Fig. 3(b), when $r^{2}$ is plotted against the rescaled input standard derivation $\sigma_{\text{in}}/\sigma_{\xi}^{1/3}$, the curves corresponding to various $\sigma_{\xi}$ collapse onto a single master curve. We thus expect a scaling law for the critical input standard deviation $\sigma_{\text{in},0}$ at which the linear-to-nonlinear transition occurs,
@@
To verify the scaling law in Eq. (13), we examine the behavior of $\sigma_{\text{in},0}$ as a function of the noise intensity $\sigma_{\xi}$. We define the transition point $\sigma_{\text{in},0}$ as the value of $\sigma_{\text{in}}$ at which the derivative $dr^{2}/d\sigma_{\text{in}}$ is maximized,
$\sigma_{\text{in},0}=\mathop{\rm argmax}\limits_{\sigma_{\text{in}}}\frac{dr^{2}(\sigma_{\text{in}})}{d\sigma_{\text{in}}}.$ (14)
To reduce the fluctuations in $\sigma_{\text{in},0}$, we fit $r^{2}$ as a function of $\sigma_{\text{in}}$ using a logarithmic logistic function of the form,
$f(x)=\frac{a}{1+e^{-b(\ln x-c)}},$ (15)
with fitting parameters $a$, $b$, and $c$ [see Fig. 3(c)]. We then obtain $\sigma_{\text{in},0}$ directly from these fitted parameters:
$\sigma_{\text{in},0}=e^{c}\left(\frac{b-1}{b+1}\right)^{1/b}.$ (16)
Figure 3(d) shows $\sigma_{\text{in},0}$ as a function of the noise intensity $\sigma_{\xi}$ for different spectral radii. The dashed lines show power-law fits of the form $\sigma_{\text{in},0}\propto\sigma_{\xi}^{\eta}$, with $\eta=0.340$ for $\rho=0.4$, $\eta=0.334$ for $\rho=0.6$, and $\eta=0.333$ for $\rho=0.8$, confirming the scaling law in Eq. (13).
@@
The scaling exponent $\eta=1/3$ might be understood from the analytic properties of the activation function $h(x)$ in Eq. (6). When both $\sigma_{\text{in}}$ and $\sigma_{\xi}$ are small, the network dynamics are primarily governed by the behavior of $h(x)$ near $x=0$. For example, the commonly used sigmoid-type activation function $h(x)=\tanh(x)$ admits a Taylor expansion of the form $h(x)=x+O(x^{3})$ around the origin. We conjecture that the observed exponent $\eta=1/3$ arises from the cubic term $O(x^{3})$, which represents the leading-order nonlinearity beyond the linear regime. To test this conjecture, we examine in Appendix E the scaling behavior with respect to the noise intensity using alternative activation functions. For an activation function whose Taylor expansion begins at fifth order, i.e., $h(x)=x+O(x^{5})$, we find that the scaling exponent becomes $\eta=1/5$ [see Figs. 9(c) and (d)]. In contrast, for a piecewise linear activation function that behaves exactly as $h(x)=x$ within the range $|x|<1$, the transition point $\sigma_{\text{in},0}$ remains constant regardless of the noise intensity, yielding $\eta=0$ [see Fig. 9(e)]. These results support the conjecture that the scaling exponent $\eta$ is determined by the inverse of the order of the lowest-order nonlinear term in the expansion of $h(x)$ around $x=0$.
@@
where $\beta$ is the ridge parameter. As discussed in Appendix C, $\beta$ corresponds to an effective noise standard deviation $\sigma_{\xi}\sim\beta^{1/2}$. Consequently, in the ridge regression case, the appropriate rescaling of the input standard deviation is given by $\sigma_{\text{in}}/\beta^{1/6}$, as demonstrated in Fig. 7.
@@ IV.3 Scaling for the network size
Figure 4: Network size scaling of the transition point. (a) Information processing capacity $r^{2}$ versus the rescaled input standard deviation $\tilde{\sigma}_{\text{in}}=\sigma_{\text{in}}/\sigma_{\xi}^{1/3}$ for $\sigma_{\xi}=10^{-8}$ at several network sizes NN (XOR task with delay $k=20$). The dashed curves show logarithmic-logistic fits. The peak value of $r^{2}$ saturates to 11 for $N\gtrsim 2000$, beyond which both $\Delta r^{2}_{\text{max}}$ and $\tilde{\sigma}_{\text{in},0}$ exhibit power-law scaling with respect to $N$. (b) Network size dependence of the maximum derivative $\Delta r^{2}_{\text{max}}$ at the transition point in a double-log plot. The dashed lines represent $\Delta r^{2}_{\text{max}}\propto N^{\alpha}$. (c) Network size dependence of the transition point $\tilde{\sigma}_{\text{in},0}$ in a double-log plot. The dashed lines represent $\tilde{\sigma}_{\text{in},0}\propto N^{-\beta}$. (d) Exponents $\alpha$ and $\beta$ for different delay values $k$. The power-law fits for $\tilde{\sigma}_{\text{in},0}$ and $\Delta r^{2}_{\text{max}}$ are performed for $N\geq 2200$. The error bars indicate the standard errors from the fitting. The fact that $\alpha>\beta$ implies that the transition sharpens faster than $\tilde{\sigma}_{\text{in},0}$ vanishes.
@@
As in Eq. (14), we define the rescaled critical point $\tilde{\sigma}_{\text{in},0}$ as the value of $\tilde{\sigma}_{\text{in}}$ at which the derivative $dr^{2}/d\tilde{\sigma}_{\text{in}}$ is maximized. In addition, we denote the maximum of the derivative by $\Delta r^{2}_{\text{max}}$:
$\Delta r^{2}_{\text{max}}=\frac{dr^{2}(\tilde{\sigma}_{\text{in},0})}{d\tilde{\sigma}_{\text{in}}}.$ (19)
@@
To reduce the fluctuations in $\tilde{\sigma}_{\text{in},0}$ and $\Delta r^{2}_{\text{max}}$, we fit $r^{2}$ as a function of $\tilde{\sigma}_{\text{in}}$ using a logarithmic logistic function given by Eq. (15) with fitting parameters $a$, $b$, and $c$. In addition to Eq. (16), $\Delta r^{2}_{\text{max}}$ is given by
@@
Figure 4(a) presents examples of fits of $r^{2}$ versus the rescaled input standard deviation $\tilde{\sigma}_{\text{in}}$ for the XOR task with delay $k=20$. The noise intensity is $\sigma_{\xi}=10^{-8}$, and the spectral radius of $W$ is $\rho=0.8$. The results show that the peak value of $r^{2}$ saturates to unity for $N\gtrsim 2000$. Figure 4(b) shows the size dependence of $\Delta r^{2}_{\text{max}}$ in a double-log plot for different delay values $k$ in the XOR task. For sufficiently large $N$, we find that the sharpness of the transition scales as $\Delta r^{2}_{\text{max}}\propto N^{\alpha}$ with a positive exponent $\alpha$. Figure 4(c) shows the location of the transition $\tilde{\sigma}_{\text{in},0}$ as a function of $N$, which scales as $\tilde{\sigma}_{\text{in},0}\propto N^{-\beta}$ with a positive exponent $\beta$. Note that both $\Delta r^{2}_{\text{max}}$ and $\tilde{\sigma}_{\text{in},0}$ enter the power-law regime once the peak value of $r^{2}$ versus $\tilde{\sigma}_{\text{in}}$ saturates to unity.
Figure 4(d) shows the exponents $\alpha$ and $\beta$ as functions of $k$. The power-law fits for $\tilde{\sigma}_{\text{in},0}$ and $\Delta r^{2}_{\text{max}}$ are performed over the range $N\in[2200,6000]$. As $k$ increases, the XOR task demands longer memory, making it more difficult. We observe that both $\alpha$ and $\beta$ grow with $k$, although we cannot rule out the possibility that this trend is influenced by the choice of fitting range. Appendix B provides a detailed discussion of how the fitting range affects the extracted values of $\alpha$ and $\beta$. We find that $\alpha>\beta$, which implies that the sharpness of the transition increases more rapidly than the decrease in $\tilde{\sigma}_{\text{in},0}$. This result suggests that, when plotting $r^{2}$ against $\tilde{\sigma}_{\text{in}}N^{\beta}$, a discontinuous transition may emerge in the limit $N\to\infty$, as illustrated in Fig. 1.